OpenStudy (ijustneedsomehelppleazzzzzzzzz):

(1 pt) Which number is irrational? A. square root of 64 B. square root of 256 C. square root of 100 D. square root of 80

Mehek (mehek14):

Any idea? :)

OpenStudy (ijustneedsomehelppleazzzzzzzzz):

b

Atsie (atsie):

Bear in mind that an irrational number is a number that cannot be written as a simple fraction most times. Why do you guess B anyhow?

Mehek (mehek14):

\(\sqrt{256}=16\) so no

OpenStudy (ijustneedsomehelppleazzzzzzzzz):

idk

OpenStudy (ijustneedsomehelppleazzzzzzzzz):

is it a

Mehek (mehek14):

\(\sqrt{64}=8\) no

OpenStudy (ijustneedsomehelppleazzzzzzzzz):

how do you do that?

Mehek (mehek14):

Because 8 * 8 = 64

OpenStudy (ijustneedsomehelppleazzzzzzzzz):

I got it

OpenStudy (ijustneedsomehelppleazzzzzzzzz):

so its d

Mehek (mehek14):

Yes